LucyReed:

What is 139 x 294 = ?

Rylee88:

If you click the three lines at the top left of your screen, You scroll down and there is a calculator for you/anyone to use. Especially for simple operations.

LucyReed:

:D

Rylee88:

1 attachment
surjithayer:

\[139\times294=139\times(300-6)\]\[=139\times300-139\times6\] \[=41700-834\]\[=40866\]

Aspect:

@mrwhskers Try not to give direct answers, it is strictly prohibited and against the rules. Explain with your answer like surjithayer. Thank you! Also use your resources like @rylee88 said.

mrwhskers:

@aspect wrote:
@mrwhskers Try not to give direct answers, it is strictly prohibited and against the rules. Explain with your answer like surjithayer. Thank you! Also use your resources like @rylee88 said.
oh im sorry. ill delete my post

Aspect:

@mrwhskers wrote:
@aspect wrote:
@mrwhskers Try not to give direct answers, it is strictly prohibited and against the rules. Explain with your answer like surjithayer. Thank you! Also use your resources like @rylee88 said.
oh im sorry. ill delete my post
It's okay thank you for complying (:

mrwhskers:

@aspect wrote:
@mrwhskers wrote:
@aspect wrote:
@mrwhskers Try not to give direct answers, it is strictly prohibited and against the rules. Explain with your answer like surjithayer. Thank you! Also use your resources like @rylee88 said.
oh im sorry. ill delete my post
It's okay thank you for complying (:
yw